There are four 3-digit natural numbers, each of them equals the sum of the
cube of its digits, three of them are:
153=1+125+27
370=27+343+0
407=64+0+343
What is the 4th one?

Answer by Edwin McCravy(20086)

There are four 3-digit natural numbers, each of them equals the sum of the
cube of its digits, three of them are:
153=1+125+27
370=27+343+0
407=64+0+343
What is the 4th one?

Since the second one has a 0 on the end, 
and since 1 is 13, just add 1
to both sides of it and get

371=27+343+1

Edwin